If P(B) = 0.4, P(A or B) = 0.7, and P(A and B) = 0.3, find P(A).

Answer by Boreal(15235)

p(a)+p(b)-p(a and b)=p(a or b)
p(a)+0.4-0.7=0.3
p(a)=0.6